# http://www.rolfdrechsler.de

### v1

@@ 1 Introduction
Test and verification are central tasks in modern system design. With the continuous increase in system complexity and size, these tasks are becoming even more critical to ensure functional correctness and reliability. In the software domain, Test Driven Development (TDD) is a well-established methodology that has demonstrated several advantages in improving quality and reducing development effort [1, 2, 3]. As a further evolution of this idea, Behavior-Driven Development (BDD) has been introduced, offering a more structured and behavior-oriented view on specification and testing [3, 4, 5, 6].
@@ 2 Scenario Generation
A central element of Behavior-Driven Development (BDD) is the definition of scenarios that describe the expected system behavior in a structured and executable form. In practice, these scenarios are commonly written using the Gherkin language. A typical Gherkin feature file may look as follows:
@@
Such scenarios must traditionally be crafted manually, a process that is both time-consuming and prone to inconsistencies or omissions, as noted in the introduction. Earlier attempts to automate scenario derivation relied on classical AI and Natural Language Processing (NLP) techniques [10]. However, these methods were limited in their ability to extract meaningful behavioral descriptions from non-trivial specifications and required substantial fine-tuning to achieve acceptable results.
@@ 3 Case study: ALU
To demonstrate the complete flow of our approach from textual specification to executable verification, we consider a 16-bit ALU. First, only an informal textual description of the intended behavior of the ALU’s is available. This mirrors common industrial scenarios, where specifications are written in natural language and must be manually translated into hardware and corresponding test scenarios. Using our LLM-based methodology, both the Verilog implementation and the associated BDD scenarios are automatically generated.
@@ 3.2 Automatic Scenario Generation
In the next step, BDD scenarios are automatically derived from the same textual specification, ensuring consistency between the implementation and verification intent. These scenarios follow the Gherkin Given-When-Then structure and describe typical as well as boundary-case behaviors. A representative examples is shown in Figure 1, where the scenario is described followed by several examples that can be directly used as test cases (see below).

### v2

@@ 1 Introduction
Test and verification are central tasks in modern system design. With the continuous increase in system complexity and size, these tasks are becoming even more critical to ensure functional correctness and reliability. In the software domain, Test Driven Development (TDD) is a well-established methodology that has demonstrated several advantages in improving quality and reducing development effort [1, 2, 3]. As a further evolution of this idea, Behavior-Driven Development (BDD) has been introduced, offering a more structured and behavior-oriented view on specification and testing [3, 4, 5, 6, 7, 8].
@@ 2 Scenario Generation
A central element of BDD is the definition of scenarios that describe the expected system behavior in a structured and executable form. In practice, these scenarios are commonly written using the Gherkin language. A typical Gherkin feature file may look as follows:
@@
Such scenarios must traditionally be crafted manually, a process that is both time-consuming and prone to inconsistencies or omissions, as noted in the introduction. Earlier attempts to automate scenario derivation relied on classical AI and NLP techniques [12]. However, these methods were limited in their ability to extract meaningful behavioral descriptions from non-trivial specifications and required substantial fine-tuning to achieve acceptable results.
@@ 3 Case study: ALU
To demonstrate the complete flow of our approach from textual specification to executable verification, we consider a 16-bit ALU. First, only an informal textual description of the intended behavior of the ALU is available. This mirrors common industrial scenarios, where specifications are written in natural language and must be manually translated into hardware and corresponding test scenarios. Using our LLM-based methodology, both the Verilog implementation and the associated BDD scenarios are automatically generated.
@@ 3.2 Automatic Scenario Generation
In the next step, BDD scenarios are automatically derived from the same textual specification, ensuring consistency between the implementation and verification intent. These scenarios follow the Gherkin Given-When-Then structure and describe typical as well as boundary-case behaviors. A representative example is shown in Figure 1, where the scenario is described followed by several examples that can be directly used as test cases (see below).